OpenStudy (anonymous):

I just need my answer confirmed: Find the derivative of f(x) = -4x2 + 11x at x = 10. =-69

OpenStudy (primeralph):

Show your steps.

OpenStudy (anonymous):

You are correct. The derivative is -8x + 11 -8(10) +11 = -69 =D

OpenStudy (anonymous):

I plugged it into my calculator.

OpenStudy (primeralph):

@Loveiskey18 Won't learn that way.

OpenStudy (anonymous):

@piglet9 it's 11x, not 11.

OpenStudy (anonymous):

-8(10) = -80 11(10) = 110 -80+110 = 30

OpenStudy (anonymous):

@primeralph ?

OpenStudy (anonymous):

@Loveiskey18 the derivative is 11 not 11x.....

OpenStudy (anonymous):

Oh okay, makes sense.